# SPHERE: Unveiling Spatial Blind Spots in Vision-Language Models Through Hierarchical Evaluation

### v3

@@ Abstract
Wenyu Zhang1, Wei En Ng2, Lixin Ma3*, Yuwen Wang2*, Jungqi Zhao4*, Allison Koenecke5 Boyang Li4, Lu Wang1 1Institute for Infocomm Research (IR), Agency for Science, Technology and Research (A*STAR) 2National University of Singapore (NUS), 3Tongji Unversity 4 Nanyang Technological University (NTU), 5 Cornell University
@@ 1 Introduction
Spatial perception and reasoning play an essential role in how vision-language models (VLMs) understand complex, context-rich environments Du et al. (2024); Liu et al. (2023a); Zeng et al. (2024). These capabilties are crucial for the deployment of VLMs in physical-world applications such as robotics, embodied AI, or human-assistive systems (Nair et al., 2022; Brohan et al., 2023; Stone et al., 2023; Leal et al., 2023).
@@
Model Single-skill Multi-skill Reasoning Overall Human baseline 95.4 92.5 89.0 93.0 Random baseline∗ 50.0 44.3 50.0 49.1 Best model score 78.2 58.6 64.7 67.9 Rank 1 Gemini 2.0 Flash GPT-4o GPT-4o GPT-4o Rank 2 GPT-4o LLaVA-OneVision SpatialBot-RGB Gemini 2.0 Flash Rank 3 LLaVA-OneVision Gemini 2.0 Flash Llama-3.2-Vision LLaVA-OneVision
@@
We evaluate several state-of-the-art VLMs and reveal significant shortcomings in their capabilities. These models especially lack the ability to understand distance and proximity, to reason from both allocentric and egocentric viewpoints, and to perform complex reasoning in a physical world context.
@@ 2.1 Spatial Capabilities in Multimodal LLMs
The majority of literature on multimodal LLMs focus on their conversational ability to respond fluently to the queries of human users through natural language (Dai et al., 2023; Liu et al., 2023b, 2024; Lyu et al., 2023) or other modalities such as audio and image Wu et al. (2023); Tang et al. (2024). More recently, works have started to explore the instillation of spatial capabilities in LLMs to better understand the physical world. Zheng et al. (2024) developed BAT based on text and audio by synthesizing data with a spatial audio simulator, and Devnani et al. (2024) trained a spatially-aware audio and text embedding model. SpatialBot Cai et al. (2024) directly includes pairs of RGB and depth images in existing datasets for training. A popular approach for vision-language models is to use existing algorithms and models (e.g. object detection, segmentation, depth estimation) to extract additional information from available data, and then incorporate these enhanced data for training or finetuning (Chen et al., 2024a; Zhao et al., 2023; AI, 2024a; Cai et al., 2024; Cheng et al., 2024). The resulting models have improved capabilities to infer spatial relationships between objects and to estimate object dimensions from natural images.
@@ 2.2 Benchmark Evaluations
Spatial reasoning is necessary for VLMs to understand and interact with the physical world. This has spurred the development of numerous benchmarks that vary in scope, tasks, and the types of spatial skills they focus on. Some benchmarks Li et al. (2024b); Yu et al. (2024) involve spatial capabilities in only a small subset of tasks. EmbSpatial-Bench Du et al. (2024) evaluates only a limited number of positional relationships, namely ‘above’, ‘below’, ‘left’, ‘right’, ‘far’ and ‘close‘. Zeng et al. (2024) tests the effects of spatial relations on visual grounding tasks. VSR Liu et al. (2023a) includes proximity and topological concepts. SpatialBench Cai et al. (2024) consists a small dataset of 120 images to evaluate spatial relationships, counting and size comparison between objects. Q-Spatial Bench Liao et al. (2024) and SpatialRGPT-Bench Cheng et al. (2024) focuses on model ability to recognize quantitative information such as metric distances and object sizes. These evaluations and datasets do not disentangle the effects of multiple spatial and/or visual skills. CLEVR Johnson et al. (2016) tests compositional reasoning with multiple spatial and visual skills, but the dataset is synthetic and the objects comprise only simple 3D shapes such as spheres, cubes and cylinders. Kamath et al. (2023) isolates the effect of spatial attributes by varying the spatial relations of objects while fixing object identities, but the collected dataset focuses only on positional relationships. We study the interaction effect of a wider range of skills, and introduce a novel study of model ability to infer about scenes or objects not explicitly observed but phyiscally plausible in the input images.
3 Proposed SPHERE
3.1 Benchmark Framework
@@ 3.2.1 Dataset Design and Statistics
Position + Counting: Questions require the model to count the number of specified objects that has a specified spatial relationship with another object in the image. We include both egocentric and allocentric viewpoints. Example questions are shown in Figure 4.
Distance + Counting: Questions require the model to count the number of specified objects that is at a specified proximity to another object in the image or the observer. An example of a question in Figure 3 is “How many crows are on the railing farther from the viewer?"
@@
Reasoning tasks: We design 2 tasks to evaluate the model’s ability to logically reason about the 3D physical world based on its understanding of the input 2D image. In addition to the final reasoning-based questions, we include understanding-based questions as an intermediate step to check the model’s comprehension of the 2D image.
Object occlusion: This task evaluates a model’s ability to infer the presence or properties of objects that may be fully occluded from the perspective of observer. From Figure 3, an example of an intermediate understanding-based question is “Which object is thicker? Fire hydrant or tree trunk?" The final reasoning-based question is “Where can a child be hiding? Behind the fire hydrant or behind the tree?" To answer this question correctly, the model needs to reason that a child is more likely to be obscured by the thicker tree.
Object manipulation: This task tests a model’s ability to reason about potential or hypothetical positions that objects can move to or be moved to within a scene, based on physical constraints. From Figure 3, an example of an intermediate understanding-based question is “What is behind the silver laptop? Empty space or window bars?" The final reasoning-based question is “Sitting at the desk, how can you move the silver laptop? Away from you or toward you?" To answer this question correctly, the model needs to reason that a person at the desk cannot move the laptop further back as the laptop is already positioned against the window bars.
@@
Model Single-skill Multi-skill Reasoning Pos. Count. Dist. Size Avg P + C D + C D + S Avg Occl. Manip. Avg Overall Human baseline 90.4 98.0 98.0 95.0 95.4 95.6 91.0 91.0 92.5 87.0 91.0 89.0 93.0 Random baseline∗ 50.0 - 50.0 50.0 50.0 - - 44.3 44.3 50.0 50.0 50.0 49.1 General purpose VLMs Phi-3.5-Vision (4B) 61.8 58.7 56.7 76.0 63.3 44.4 36.1 47.2 42.6 54.4 55.5 54.9 54.3 LLaVA-NeXT (7B) 53.6 68.7 54.5 70.9 61.9 42.0 35.4 36.1 37.8 48.7 49.1 48.9 50.6 LLaVA-OneVision (7B) 60.2 76.1 64.2 84.0 71.1 53.3 53.8 52.9 53.3 53.0 59.6 56.3 61.5 Qwen2-VL (7B) 63.2 74.1 53.0 78.6 67.2 36.1 33.5 50.8 40.1 54.0 55.2 54.6 55.0 InstructBLIP (8B) 45.6 64.7 50.4 56.2 54.2 32.0 28.5 38.1 32.8 54.8 52.7 53.7 47.0 Idefics2 (8B) 50.1 45.3 49.1 58.5 50.7 22.5 27.8 40.1 30.1 51.0 53.3 52.1 44.2 InternVL2.5 (8B) 62.2 72.6 64.6 78.3 69.4 50.2 43.8 41.8 45.3 52.7 53.6 53.1 57.3 Qwen-VL (10B) 55.9 72.9 59.2 72.0 65.0 37.2 29.9 34.7 33.9 54.3 54.3 54.3 52.0 Llama-3.2-Vision (11B) 58.4 53.5 52.9 67.2 58.0 29.8 27.7 40.8 32.8 60.7 56.1 58.4 49.7 Qwen2-VL (72B) 62.8 80.6 60.8 85.5 72.4 37.3 38.0 69.3 48.2 53.9 55.0 54.4 59.8 Llama-3.2-Vision (90B) 66.0 59.2 65.2 75.8 66.5 40.1 31.8 52.8 41.6 61.6 55.3 58.4 56.2 Gemini 2.0 Flash 68.6 82.1 75.2 86.9 78.2 50.3 52.5 49.7 50.9 48.0 49.2 48.6 61.7 GPT-4o 71.7 76.6 71.3 89.4 77.3 54.4 62.7 58.8 58.6 66.0 63.3 64.7 67.9 Spatial VLMs SpatialBot-RGB (3B) 55.9 67.7 52.8 73.0 62.3 38.5 39.2 38.7 38.8 57.6 60.3 59.0 53.6 SpatialBot (3B) 55.5 72.1 51.9 75.4 63.7 39.1 36.1 38.9 38.0 56.5 54.6 55.5 53.1 SpaceMantis (8B) 52.7 52.7 57.0 63.1 56.4 28.4 43.7 45.6 39.2 56.6 54.2 55.4 50.4 SpatialRGPT-RGB (8B) 59.3 70.1 59.2 74.6 65.8 42.6 46.2 40.9 43.2 54.5 53.2 53.8 55.3
@@
Model Position Position + Counting Allo. Ego. Allo. Ego. General purpose VLMs Phi-3.5-Vision (4B) 77.9 44.5 53.3 29.7 LLaVA-NeXT (7B) 69.3 36.7 47.6 32.8 LLaVA-OneVision (7B) 73.7 45.7 62.9 37.5 Qwen2-VL (7B) 69.1 56.0 37.1 37.5 InstructBLIP (8B) 43.0 48.4 31.4 32.8 Idefics2 (8B) 48.5 51.9 20.0 26.6 InternVL2.5 (8B) 75.7 47.7 59.8 34.4 Qwen-VL (10B) 65.6 45.5 41.5 30.0 Llama-3.2-Vision (11B) 60.5 56.0 31.6 26.9 Qwen2-VL (72B) 70.3 55.6 36.2 35.9 Llama-3.2-Vision (90B) 72.9 58.6 41.1 38.4 Gemini 2.0 Flash 83.2 52.9 55.2 42.2 GPT-4o 79.5 63.4 62.9 40.6 Spatial VLMs SpatialBot-RGB (3B) 68.6 42.2 47.6 23.4 SpatialBot (3B) 69.7 40.2 46.7 26.6 SpaceMantis (8B) 56.2 48.8 31.4 23.4 SpatialRGPT-RGB (8B) 73.2 44.4 48.6 32.8
@@
Model Object occlusion Object manipulation Intermed. Final Final (w/ gt intermed.) Intermed. Final Final (w/ gt intermed.) General purpose VLMs Phi-3.5-Vision (4B) 60.6 54.4 67.0 59.8 55.5 57.6 LLaVA-NeXT (7B) 60.4 48.7 54.4 58.5 49.1 48.8 LLaVA-OneVision (7B) 63.9 53.0 59.8 64.2 59.6 57.1 Qwen2-VL (7B) 63.4 54.0 60.8 55.1 55.2 55.4 InstructBLIP (8B) 48.7 54.8 57.8 52.9 52.7 45.7 Idefics2 (8B) 56.5 51.0 60.0 51.1 53.3 60.7 InternVL2.5 (8B) 63.7 52.7 58.9 62.9 53.6 57.2 Qwen-VL (10B) 57.8 54.3 57.3 54.5 54.3 49.0 Llama-3.2-Vision (11B) 52.3 60.7 62.5 53.7 56.1 57.5 Qwen2-VL (72B) 65.6 53.9 62.4 68.5 55.0 76.9 Llama-3.2-Vision (90B) 62.7 61.6 68.0 59.8 55.3 74.0 Gemini 2.0 Flash 62.4 48.0 61.0 65.5 49.2 58.3 GPT-4o 67.3 66.0 72.0 68.5 63.3 76.4 Spatial VLMs SpatialBot-RGB (3B) 59.9 57.6 65.9 60.8 60.3 54.0 SpatialBot (3B) 58.6 56.5 63.2 59.2 54.6 56.1 SpaceMantis (8B) 63.1 56.6 62.4 60.1 54.2 57.6 SpatialRGPT-RGB (8B) 61.1 54.5 63.2 60.1 53.2 58.5

### v4

@@ Abstract
Wenyu Zhang1, Wei En Ng2, Lixin Ma3*, Yuwen Wang2*, Junqi Zhao4*, Allison Koenecke5 Boyang Li4, Lu Wang1 1Institute for Infocomm Research (IR), Agency for Science, Technology and Research (A*STAR) 2National University of Singapore (NUS), 3Tongji Unversity 4Nanyang Technological University (NTU), 5Cornell University
@@ 1 Introduction
Spatial perception and reasoning play an essential role in how vision-language models (VLMs) understand complex, context-rich environments Du et al. (2024); Liu et al. (2023a); Zeng et al. (2024). These capabilities are crucial for the deployment of VLMs in physical-world applications such as robotics, embodied AI, or human-assistive systems (Nair et al., 2022; Brohan et al., 2023; Stone et al., 2023; Leal et al., 2023).
@@
Model Single-skill Multi-skill Reasoning Overall Human baseline 95.4 92.5 89.0 93.0 Random baseline∗ 50.0 44.3 50.0 49.1 Best model score 78.2 58.6 64.7 67.9 Rank 1 Gemini 2.0 Flash GPT-4o GPT-4o GPT-4o Rank 2 GPT-4o Qwen2.5-VL SpatialBot-RGB Qwen2.5-VL Rank 3 Qwen2.5-VL LLaVA-OneVision Llama-3.2-Vision Gemini 2.0 Flash
@@
We evaluate state-of-the-art VLMs and reveal significant shortcomings in their capabilities. These models especially lack the ability to understand distance and proximity, to reason from both allocentric and egocentric viewpoints, and to perform complex reasoning in a physical world context.
@@ 2.1 Spatial Capabilities in Multimodal LLMs
The majority of literature on multimodal LLMs focus on their conversational ability to respond fluently to the queries of human users through natural language (Dai et al., 2023; Liu et al., 2023b, 2024; Lyu et al., 2023) or other modalities such as audio and image Wu et al. (2023); Tang et al. (2024). More recently, works have started to explore the integration of spatial capabilities in LLMs to better understand the physical world. Zheng et al. (2024) developed BAT based on text and audio by synthesizing data with a spatial audio simulator, and Devnani et al. (2024) trained a spatially-aware audio and text embedding model. SpatialBot Cai et al. (2024) directly includes pairs of RGB and depth images in existing datasets for training. A popular approach for vision-language models is to use existing algorithms and models (e.g. object detection, segmentation, depth estimation) to extract additional information from available data, and then incorporate these enhanced data for training or finetuning (Chen et al., 2024a; Zhao et al., 2023; AI, 2024a; Cai et al., 2024; Cheng et al., 2024). The resulting models have improved capabilities to infer spatial relationships between objects and to estimate object dimensions from natural images.
@@ 2.2 Benchmark Evaluations
Spatial reasoning is necessary for VLMs to understand and interact with the physical world. This has spurred the development of numerous benchmarks that vary in scope, tasks, and the types of spatial skills they focus on. Some benchmarks Li et al. (2024b); Yu et al. (2024) involve spatial capabilities in only a small subset of tasks. EmbSpatial-Bench Du et al. (2024) evaluates only a limited number of positional relationships, namely ‘above’, ‘below’, ‘left’, ‘right’, ‘far’ and ‘close‘. Zeng et al. (2024) tests the effects of spatial relations on visual grounding tasks. VSR Liu et al. (2023a) includes proximity and topological concepts. SpatialBench Cai et al. (2024) consists of a small dataset of 120 images to evaluate spatial relationships, counting and size comparison between objects. Q-Spatial Bench Liao et al. (2024) and SpatialRGPT-Bench Cheng et al. (2024) focuses on model ability to recognize quantitative information such as metric distances and object sizes. These evaluations and datasets do not disentangle the effects of multiple spatial and/or visual skills. CLEVR Johnson et al. (2016) tests compositional reasoning with multiple spatial and visual skills, but the dataset is synthetic and the objects comprise only simple 3D shapes such as spheres, cubes and cylinders. Kamath et al. (2023) isolates the effect of spatial attributes by varying the spatial relations of objects while fixing object identities, but the collected dataset focuses only on positional relationships. We study the interaction effect of a wider range of skills, and introduce a novel study of model ability to infer about scenes or objects not explicitly observed but physically plausible in the input images.
3 The SPHERE Benchmark
3.1 Overview
@@ 3.2.1 Dataset Design and Statistics
Position + Counting: Questions require the model to count the number of specified objects that have a specified spatial relationship with another object in the image. We include both egocentric and allocentric viewpoints. Example questions are shown in Figure 4.
Distance + Counting: Questions require the model to count the number of specified objects that are at a specified proximity to another object in the image or the observer. An example of a question in Figure 3 is “How many crows are on the railing farther from the viewer?”
@@
Reasoning tasks: We design 2 tasks to evaluate the model’s ability to logically reason about the 3D physical world based on its understanding of the input 2D image. In addition to the final reasoning-based questions, we include understanding-based questions as an intermediate step to check the model’s comprehension of the 2D image. In our design, “intermediate” questions focus on assessing the perception and understanding of visible objects and scenes. These tasks align with cognitive theories like Piaget’s stages of cognitive development, which suggest that perceptual skills are crucial for building more abstract reasoning abilities assessed through the “final” reasoning questions. The “final” reasoning questions require the ability to infer the existence, position, or state of partially or non-visible objects and scenes that are physically plausible. These inferences rely on understanding spatial relationships and object permanence, as the missing information could only be revealed by changing viewpoints or interacting with the scene, such as moving objects to uncover hidden ones.
Object occlusion: This task evaluates a model’s ability to infer the presence or properties of objects that may be fully occluded from the perspective of observer. From Figure 3, an example of an intermediate understanding-based question is “Which object is thicker? Fire hydrant or tree trunk?” The final reasoning-based question is “Where can a child be hiding? Behind the fire hydrant or behind the tree?” To answer this question correctly, the model needs to reason that a child is more likely to be obscured by the thicker tree.
Object manipulation: This task tests a model’s ability to reason about potential or hypothetical positions that objects can move to or be moved to within a scene, based on physical constraints. From Figure 3, an example of an intermediate understanding-based question is “What is behind the silver laptop? Empty space or window bars?” The final reasoning-based question is “Sitting at the desk, how can you move the silver laptop? Away from you or toward you?” To answer this question correctly, the model needs to reason that a person at the desk cannot move the laptop further back as the laptop is already positioned against the window bars.
@@
Model Single-skill Multi-skill Reasoning Pos. Count. Dist. Size Avg P + C D + C D + S Avg Occl. Manip. Avg Overall Human baseline 90.4 98.0 98.0 95.0 95.4 95.6 91.0 91.0 92.5 87.0 91.0 89.0 93.0 Random baseline∗ 50.0 - 50.0 50.0 50.0 - - 44.3 44.3 50.0 50.0 50.0 49.1 General purpose VLMs Phi-3.5-Vision (4B) 61.8 58.7 56.7 76.0 63.3 44.4 36.1 47.2 42.6 54.4 55.5 54.9 54.3 LLaVA-NeXT (7B) 53.6 68.7 54.5 70.9 61.9 42.0 35.4 36.1 37.8 48.7 49.1 48.9 50.6 LLaVA-OneVision (7B) 60.2 76.1 64.2 84.0 71.1 53.3 53.8 52.9 53.3 53.0 59.6 56.3 61.5 Qwen2-VL (7B) 63.2 74.1 53.0 78.6 67.2 36.1 33.5 50.8 40.1 54.0 55.2 54.6 55.0 Qwen2.5-VL (7B) 59.3 82.1 58.9 73.6 68.5 44.7 31.6 37.9 38.1 48.7 54.1 51.4 54.1 Janus-Pro (7B) 57.1 70.6 59.3 76.5 65.9 37.3 33.5 46.8 39.2 53.5 54.3 53.9 54.0 InstructBLIP (8B) 45.6 64.7 50.4 56.2 54.2 32.0 28.5 38.1 32.8 54.8 52.7 53.7 47.0 Idefics2 (8B) 50.1 45.3 49.1 58.5 50.7 22.5 27.8 40.1 30.1 51.0 53.3 52.1 44.2 InternVL2.5 (8B) 62.2 72.6 64.6 78.3 69.4 50.2 43.8 41.8 45.3 52.7 53.6 53.1 57.3 Qwen-VL (10B) 55.9 72.9 59.2 72.0 65.0 37.2 29.9 34.7 33.9 54.3 54.3 54.3 52.0 Llama-3.2-Vision (11B) 58.4 53.5 52.9 67.2 58.0 29.8 27.7 40.8 32.8 60.7 56.1 58.4 49.7 Qwen2-VL (72B) 62.8 80.6 60.8 85.5 72.4 37.3 38.0 69.3 48.2 53.9 55.0 54.4 59.8 Qwen2.5-VL (72B) 63.7 81.1 69.9 89.3 76.0 51.5 55.7 66.4 57.9 50.3 56.4 53.3 64.3 Llama-3.2-Vision (90B) 66.0 59.2 65.2 75.8 66.5 40.1 31.8 52.8 41.6 61.6 55.3 58.4 56.2 Gemini 2.0 Flash 68.6 82.1 75.2 86.9 78.2 50.3 52.5 49.7 50.9 48.0 49.2 48.6 61.7 GPT-4o 71.7 76.6 71.3 89.4 77.3 54.4 62.7 58.8 58.6 66.0 63.3 64.7 67.9 Spatial VLMs SpatialBot-RGB (3B) 55.9 67.7 52.8 73.0 62.3 38.5 39.2 38.7 38.8 57.6 60.3 59.0 53.6 SpatialBot (3B) 55.5 72.1 51.9 75.4 63.7 39.1 36.1 38.9 38.0 56.5 54.6 55.5 53.1 SpaceMantis (8B) 52.7 52.7 57.0 63.1 56.4 28.4 43.7 45.6 39.2 56.6 54.2 55.4 50.4 SpatialRGPT-RGB (8B) 59.3 70.1 59.2 74.6 65.8 42.6 46.2 40.9 43.2 54.5 53.2 53.8 55.3
@@
Model Position Position + Counting Allo. Ego. Allo. Ego. General purpose VLMs Phi-3.5-Vision (4B) 77.9 44.5 53.3 29.7 LLaVA-NeXT (7B) 69.3 36.7 47.6 32.8 LLaVA-OneVision (7B) 73.7 45.7 62.9 37.5 Qwen2-VL (7B) 69.1 56.0 37.1 37.5 Qwen2.5-VL (7B) 69.0 48.8 49.1 37.5 Janus-Pro (7B) 65.0 48.6 37.1 37.5 InstructBLIP (8B) 43.0 48.4 31.4 32.8 Idefics2 (8B) 48.5 51.9 20.0 26.6 InternVL2.5 (8B) 75.7 47.7 59.8 34.4 Qwen-VL (10B) 65.6 45.5 41.5 30.0 Llama-3.2-Vision (11B) 60.5 56.0 31.6 26.9 Qwen2-VL (72B) 70.3 55.6 36.2 35.9 Qwen2.5-VL (72B) 78.7 47.6 58.1 40.6 Llama-3.2-Vision (90B) 72.9 58.6 41.1 38.4 Gemini 2.0 Flash 83.2 52.9 55.2 42.2 GPT-4o 79.5 63.4 62.9 40.6 Spatial VLMs SpatialBot-RGB (3B) 68.6 42.2 47.6 23.4 SpatialBot (3B) 69.7 40.2 46.7 26.6 SpaceMantis (8B) 56.2 48.8 31.4 23.4 SpatialRGPT-RGB (8B) 73.2 44.4 48.6 32.8
@@
Model Object occlusion Object manipulation Intermed. Final Final (w/ gt intermed.) Intermed. Final Final (w/ gt intermed.) General purpose VLMs Phi-3.5-Vision (4B) 60.6 54.4 67.0 59.8 55.5 57.6 LLaVA-NeXT (7B) 60.4 48.7 54.4 58.5 49.1 48.8 LLaVA-OneVision (7B) 63.9 53.0 59.8 64.2 59.6 57.1 Qwen2-VL (7B) 63.4 54.0 60.8 55.1 55.2 55.4 Qwen2.5-VL (7B) 65.2 48.7 51.1 61.1 54.1 55.4 Janus-Pro (7B) 58.4 53.5 55.0 62.6 54.3 53.0 InstructBLIP (8B) 48.7 54.8 57.8 52.9 52.7 45.7 Idefics2 (8B) 56.5 51.0 60.0 51.1 53.3 60.7 InternVL2.5 (8B) 63.7 52.7 58.9 62.9 53.6 57.2 Qwen-VL (10B) 57.8 54.3 57.3 54.5 54.3 49.0 Llama-3.2-Vision (11B) 52.3 60.7 62.5 53.7 56.1 57.5 Qwen2-VL (72B) 65.6 53.9 62.4 68.5 55.0 76.9 Qwen2.5-VL (72B) 67.3 50.3 57.6 64.2 56.4 66.9 Llama-3.2-Vision (90B) 62.7 61.6 68.0 59.8 55.3 74.0 Gemini 2.0 Flash 62.4 48.0 61.0 65.5 49.2 58.3 GPT-4o 67.3 66.0 72.0 68.5 63.3 76.4 Spatial VLMs SpatialBot-RGB (3B) 59.9 57.6 65.9 60.8 60.3 54.0 SpatialBot (3B) 58.6 56.5 63.2 59.2 54.6 56.1 SpaceMantis (8B) 63.1 56.6 62.4 60.1 54.2 57.6 SpatialRGPT-RGB (8B) 61.1 54.5 63.2 60.1 53.2 58.5